Activity 1: Draw Me


Directions: Given the following information, construct the rejection region. Show the solution in a step-by-step procedure.


1. H0 : = 84


Ha : 84


m= 87, s= 10, n = 35, alpha= 0.05



2. H0 : = 45


Ha : < 45


m= 40, s = 12, n = 32, alpha= 0.01


1
Expert's answer
2022-06-06T08:15:30-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu=84"

"H_1:\mu\not=84"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=34" and the critical value for a two-tailed test is "t_c = 2.032244."

The rejection region for this two-tailed test is "R = \{t:|t|> 2.032244\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{87-84}{10/\sqrt{35}}=1.7748"


Since it is observed that "|t|=1.7748< 2.032244=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed, "df=34" degrees of freedom, "t=-4" is "p=0.084885," and since "p=0.084885>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 84, at the "\alpha = 0.05" significance level.


2. The following null and alternative hypotheses need to be tested:

"H_0:\mu=45"

"H_1:\mu<45"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=31" and the critical value for a left-tailed test is "t_c = -2.452824."

The rejection region for this left-tailed test is "R = \{t:t<-2.452824\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{40-45}{12/\sqrt{32}}=-2.3570"


Since it is observed that "|t|=-2.3570> -2.452824=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for left-tailed, "df=31" degrees of freedom, "t=-2.3570" is "p=0.012459," and since "p=0.012459>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 45, at the "\alpha = 0.01" significance level.